A body of mass `m`and density `d1` lies in a non viscous liquid of density `d2`.the minimum work done to lift the body with uniform velocity through a height `h in the liquid is

Grade:11

2 Answers

Shailendra Kumar Sharma
6 years ago
So a body of mass m and density d1 to be lifted in another environment of d2 to a height h, this way the work done will be.Two force acting mg-mg(d2/d1) so work done will be MGH(1-d2/d1)
Samruddhi Khamankar
3 years ago
lifting force = apparent weight 
weight apparent = actual weight - buoyancy force 
 = mg - d1( density of liquid) *volume of liq.*gravity
[density of body (d2 )= mass of body ÷ volume of body.... therefore ,volume of body = mass of body ÷ density of body (d2) ...(1)] 
apparent weight = mg (1- d1/d2) 
work done by lifting force = mg(1- d1/d2)h ........since (W=F.S)
therefore Work done by lifting force = mgh(1-d1/d2)